OpenStudy (anonymous):

solve using the principle of zero products (s+4)(s+16)=0 choose the correct solutions for s a. 2, 4 b. -2, 4 c. -2, -4 d. -4, -16

OpenStudy (anonymous):

(s+4)=0 (s+16)=0 s = -4 s=-16 d

OpenStudy (anonymous):

can u help with this one solve using the principle of zero products: c(c+6)=0 select correct choice below : a. the solution(s) is/are c=__? - use comma to separate answers as needed type each solution only once b. there is no solution

OpenStudy (anonymous):

c(c+6)=0 c = 0 (c+6)=0 c =-6